# Reducing Energy Bloat in Large Model Training

### v2

@@ 6.2. Reducing Energy Bloat
We start with overall energy bloat reduction – intrinsic bloat without stragglers (§6.2.1) and intrinsic + extrinsic bloat with stragglers (§6.2.2) – achieved by Perseus and EnvPipe. All numbers were obtained by running on testbed GPUs, and all solutions use the same amount of GPU resources.
@@ 6.2.1. Intrinsic Bloat Reduction Without Stragglers
Table 3 compares the energy savings achieved by Perseus’s minimum iteration time energy schedule (leftmost point of the “iteration time–energy” frontier) and that by EnvPipe.
We make two observations regarding Perseus. First, models show varying amounts of energy savings because (1) their stage imbalance vary (Table 1), and (2) their forward and backward are composed on different computations, which affects time/energy sensitivity when changing the frequency.
Second, A40 demonstrates more energy savings compared to A100. This is because the dynamic clock frequency range of A100 (210–1410 MHz) is smaller than that of A40 (210–1740 MHz). Thus, tuning down the GPU’s frequency yields a relatively smaller change in computation time and energy compared to those at the maximum frequency. However, we expect the more recent GPUs to have better percentage savings due to higher maximum frequency (e.g., 1980 MHz for H100 SXM (H100whitepaper, )) and absolute savings due to high TDP (e.g., 1,200 W for each GPU on GB200 (blackwell-news, )).
EnvPipe in general provides lower energy savings, primarily due to its assumption that the final stage of a pipeline is always the heaviest, which is not always true. Additionally, it sometimes considerably degrades iteration time because it is not aware of single-choice operations inside the pipeline (§4.4) and can slow down some computations too much.
@@ 6.2.2. Intrinsic + Extrinsic Bloat Reduction With Stragglers
When stragglers create extrinsic energy bloat, the amount of energy savings for a non-straggler pipeline depends on how much energy reduction its “iteration time–energy” Pareto frontier yields for longer iteration times. Table 4 shows the amount of energy savings of a pipeline given varying degrees of straggler slowdown. As described in Section 3.1, a non-straggler pipeline’s iteration time is set to be $T_{\textrm{opt}}=\min(T^{*},T^{\prime})$, and energy reduction comes from (1) slowing down the pipeline itself and (2) reducing the time blocking on communication, waiting for the straggler. Slowing down the non-stragglers beyond $T^{*}$ will increase total energy consumption; hence, Perseus does not slow down pipelines further than $T^{*}$.
The percentage of savings first increases with the straggler’s iteration time, but then slowly wanes as it slows down beyond $T^{*}$. This is expected. The absolute amount of energy reduction in Joules is the largest when the non-straggler pipeline’s iteration time is $T^{*}$ and constant afterward, because Perseus does not slow down non-straggler pipelines beyond $T^{*}$. Thus, as the straggler slows down beyond $T^{*}$, the time and energy consumed while waiting for the straggler increases, lowering the percentage of energy savings.
@@
6.2.3. How Much Potential Savings Was Realized?
The largest possible savings under our problem setting occurs when running every computation at their minimum-energy frequencies (i.e., the $T^{*}$ point on the “iteration time–energy” Pareto frontier). For intrinsic bloat without stragglers, Perseus on average realizes 74% and 89% the opportunity on A100 and A40, respectively, with negligible slowdown. This is possible because there are much more non-critical computations in the DAG that can be slowed down than critical ones. With stragglers, Perseus fully realizes potential savings when the straggler’s slowdown degrees are on average 1.1 and 1.15 on A100 and A40 respectively, which is not unrealistic considering slowdowns reported in literature (§2.3).
@@ 6.3. Large-Scale Emulation
We profile the time and energy consumption of each layer (e.g., Transformer decoder) in GPT-3 175B and Bloom 176B and run our optimization algorithm to obtain the “iteration time–energy” frontier. We perform strong scaling when varying the number of GPUs (Table 5) in order to keep the global batch size constant (largebatch1, ; largebatch2, ). We used A100 SXM GPUs for emulation, which we believe are more representative of large-scale training infrastructure.
@@
We compare the percentage of energy savings vs. running all maximum frequencies given by our emulator and real experiments for our A100 workloads and find that the emulator always underestimates energy savings. Specifically, savings on the leftmost and rightmost point of the Pareto frontier are underestimated by 18.6% and 21.7% on average, respectively. We believe this is due to our simplifying assumption that $P_{\mathrm{blocking}}$ is constant regardless of the GPU’s frequency. This means savings given by the emulator can be considered a lower bound for actual savings.
@@
Results Summary.
Figure 7 breaks down the amount of energy bloat reduction for GPT-3 175B and Bloom 176B when slowdown degree is 1.2 on 1,024 GPUs. EnvPipe can only reduce intrinsic bloat as it does not provide an “iteration time–energy” frontier; even for intrinsic bloat, it is suboptimal. In contrast, Perseus reduces energy consumption by up to 30% by reducing both intrinsic and extrinsic energy bloat.
@@
Table 6 shows Perseus’s intrinsic energy bloat reduction without stragglers for GPT-3 175B and Bloom 176B. The number of microbatches was varied based on Table 5. For all models, as more and more microbatches are added to the pipeline, the amount of intrinsic bloat decreases. This is fundamentally due to the ratio of microbatches in the 1F1B’s warm-up and flush phase (beginning and end) vs. steady state phase (middle). Most microbatches in the warm-up and flush phases can slow down until their minimum energy frequency, yielding large energy savings. However, microbatches in the pipeline’s steady state cannot slow down to their full potential when the amount of stage imbalance is not large, thereby yielding modest savings. When the number of microbatches in the pipeline increases, only the number of steady state microbatches increases, and energy reduction converges to the average energy savings of steady state microbatches.
@@
We introduce stragglers in large-scale emulation in various strong scaling configurations (Table 5) and with varying straggler slowdown degree. Figure 8 reports the amount of intrinsic + extrinsic energy bloat reduction achieved by Perseus. The trend where energy saving increases until $T^{\prime}<T^{*}$ and wanes afterward is consistent with Section 6.2.2.
An interesting observation here is that there is a tradeoff between scale and energy savings: configurations with more pipelines have less percentage of energy savings or less amount of energy savings per pipeline. It may seem intuitive to assume that more pipelines brings more energy savings, as there is only one straggler pipeline and all the other pipelines can reduce their energy consumption. However, this holds only in weak scaling configuration, where per-pipeline batch size is held constant (i.e., increasing the global batch size proportionally to the number of pipelines). Instead, in the more realistic strong scaling configuration (i.e., the global batch size is constant and per-pipeline batch size is decreased as more pipelines deployed), each pipeline’s number of microbatches changes. With fewer microbatches, the ratio of pipeline bubble (time that GPUs are idle) at the beginning and end of each pipeline iteration increases (megatronlm-sc21, ). These bubbles cannot be perfectly eliminated by intrinsic energy bloat reduction, resulting in a smaller energy saving percentages. However, as the absolute number of GPUs increase, even a small savings percentage gives huge absolute energy savings.
@@ 6.4. Iteration Time–Energy Frontier Comparison
The energy bloat reductions in Sections 6.2 and 6.3 were made possible by the “iteration time–energy” frontier obtained using Perseus’s optimization algorithm (§4). Here, we further examine the frontier with different parallelization configurations and models and compare against Zeus (zeus-nsdi23, ). Since Zeus only produces the training time–energy frontier for single-GPU training jobs, we implemented two Zeus-based baselines for large model training scenarios:
@@
Figure 9 shows the frontiers of all solutions for different sizes of GPT-3 under three parallelization configurations: (a) four stage pipeline parallelism on A100; (b) eight stage pipeline parallelism on A40; and (c) 3D parallelism (data parallelism 2, tensor parallelism 2, pipeline parallelism 4) on A40. All results were obtained from running on testbed GPUs. Appendix G has frontiers for other models.
Perseus Pareto-dominates Zeus. ZeusGlobal is unaware of pipeline stage imbalances and slows down every stage, unable to reduce intrinsic energy bloat. ZeusPerStage can balance the forward computation time of each stage, but is unaware of the critical path of the DAG, slowing down critical computations. In contrast, Perseus can precisely slow down non-critical computations, effectively reducing energy bloat.
@@ 6.5. Overhead of Perseus
In the beginning of training, the client-side profiler (§5) profiles forward and backward computations in each stage. For our A100 workloads, the initial profiling phase increased end-to-end training time by 13 minutes on average, which is negligible overhead for large model training.
@@
The average runtime of the optimization algorithm (§4) across our A100 workloads was 6.5 minutes, with the longest being Bloom 3B (15.7 minutes). For our largest scale emulation (GPT-3 175B on A100 with 8,192 GPUs), the algorithm ran for 87 seconds, which is shorter because it is sufficient to optimize just one data parallel copy (§4.4). While runtime will increase with larger DAGs for larger models, we believe the overhead is justified because training time also increases with the scale of the training job. Looking up the optimal energy schedule given $T^{\prime}$ is instant.
7. Related Works
@@
Many recent works focus on enabling and accelerating large model training using 3D parallelism. GPipe (gpipe-neurips19, ) and PipeDream (pipedream-sosp19, ) were the first to introduce pipeline parallelism. Megatron-LM (megatronlm-arxiv, ; megatronlm-sc21, ) is a large model training framework that provides efficient 3D parallel execution plans tailored for Transformer models. 3D parallelism is considered to be go-to solutions for modern large model development due to strong open-source implementations (megatronlm-github, ; deepspeed-github, ; gpt-neox-github, ) and relatively simple implementation complexity. Finally, Alpa (alpa-osdi22, ) and GSPMD (spmd_socc22, ) provide automatic parallelization for general DNNs. However, energy consumption is not an optimization metric for any of the major large model training frameworks.
DNN Training and Energy Consumption.
A recent line of work has highlighted the enormous amount of energy consumption and carbon emission of DNN training (patterson2021carbon, ; energy-nlp-policy, ; quantify-carbon, ; dodge-faact22, ; bloomcarbon-arxiv23, ) and proposed optimization methods (odpp, ; gpoeo, ; zeus-nsdi23, ; chase-ccai23, ; depo, ; envpipe-atc23, ). Especially, Zeus (zeus-nsdi23, ) is a recent work that observes the tradeoff between GPU computation time and energy consumption, but focuses on single-GPU training. EnvPipe (envpipe-atc23, ), on the other hand, aims to reduce the energy consumption of large model training with minimum slowdown. However, its heuristic assumes that the last pipeline stage is always the bottleneck, leading to suboptimal savings. Perseus is a superset of both Zeus and EnvPipe, achieved by viewing large model training as a computation DAG and introducing a principled optimization algorithm. Perseus is also the first to introduce the notion of energy bloat.
@@ 8. Conclusion
We presented Perseus, a software-based energy optimization system for large model training. Perseus builds on the observation that there are computation imbalances at different levels in large model training that causes intrinsic and extrinsic energy bloat, and introduces a principled graph cut-based algorithm that simultaneously reduces both. As a result, Perseus advances the state-of-the-art of DNN training energy optimization by delivering energy savings with negligible slowdown, thereby reducing average power draw as well.
@@ B.1. Minimum Imbalance Pipeline Partitioning
B.2. Alternative Planning Methods

### v3

@@ 6.2. Reducing Energy Bloat
We start with overall energy bloat reduction—intrinsic bloat without stragglers (§6.2.1) and intrinsic + extrinsic bloat with stragglers (§6.2.2)—achieved by Perseus and EnvPipe. All numbers were obtained by running on testbed GPUs. All solutions use the same amount of GPU hardware resources.
@@ 6.2.1. Intrinsic Bloat Reduction Without Stragglers
Table 3 compares the energy savings achieved by Perseus’s minimum iteration time energy schedule (leftmost point of the time–energy frontier) and that by EnvPipe.
We make two observations regarding Perseus. First, models show varying amounts of energy savings because (1) their stage imbalances vary (Table 1), and (2) their forward and backward are composed of different computations, which affects time/energy sensitivity when changing the frequency. For instance, unlike other models, Wide-ResNet 1.5B on A100 after minimum imbalance stage partitioning has nearly perfect stage balance, resulting in low intrinsic energy bloat. However, as will be seen in Section 6.2.2, such models tend to achieve greater extrinsic bloat savings because most of their computations run at a high frequency, and slowing them down due to stragglers leads to higher energy reduction.
Second, A40 demonstrates more energy savings compared to A100. This is because the dynamic clock frequency range of A100 (210–1410 MHz) is smaller than that of A40 (210–1740 MHz). Thus, tuning down the GPU’s frequency yields a relatively smaller change in computation time and energy compared to those at the maximum frequency. However, we expect the more recent GPUs to have better percentage savings due to higher maximum frequency (e.g., 1980 MHz for H100 SXM (H100whitepaper, )) and better absolute savings due to high TDP (e.g., 1,200 W for each GPU on GB200 (blackwell-news, )).
EnvPipe in general provides lower energy savings, primarily due to its assumption that the final stage of a pipeline is always the heaviest. This is only correct with a probability of $1/N$, where $N$ is the number of pipeline stages. Additionally, it sometimes considerably degrades iteration time because it is not aware of single-choice operations inside the pipeline (§4.4) and can slow down some computations too much.
@@ 6.2.2. Intrinsic + Extrinsic Bloat Reduction With Stragglers
When stragglers create extrinsic energy bloat, the amount of energy savings depends on how much energy reduction the time–energy frontier yields for longer iteration times. Table 4 shows the amount of energy savings of a non-straggler pipeline given varying straggler slowdowns. For a given slowdown factor ($T^{\prime}/T$), a non-straggler pipeline’s iteration time is set to be $T_{\textrm{opt}}=\min(T^{*},T^{\prime})$3.1), and energy reduction comes from (1) slowing down the pipeline itself and (2) reducing the time (and energy) blocking on communication, waiting for the straggler.
The percentage of savings initially increases with the straggler’s iteration time, but then gradually decreases as it slows down beyond $T^{*}$. This is expected. The absolute amount of energy reduction in Joules is the largest when the straggler’s iteration time is $T^{*}$ and constant afterward, because Perseus does not slow down non-straggler pipelines beyond $T^{*}$3.1). Thus, as the straggler slows down beyond $T^{*}$, additional time and energy is consumed while waiting for the straggler, lowering the percentage of energy savings.
@@
6.2.3. How Much Potential Saving Was Realized?
The largest possible savings under our problem setting occurs when running every computation at their minimum-energy frequencies (i.e., the $T^{*}$ point on the time–energy frontier). For intrinsic bloat without stragglers, Perseus realizes on average 74% and 89% of the potential savings on A100 and A40, respectively, with negligible slowdown. This is possible because there are much more non-critical computations in the DAG that can be slowed down than critical ones. With stragglers, Perseus fully realizes potential savings when the straggler’s slowdown degrees are on average 1.1 and 1.15 on A100 and A40 respectively, which is not unrealistic considering slowdowns reported in literature (§2.3).
@@ 6.3. Large-Scale Emulation
We profile the time and energy consumption of each layer (e.g., Transformer decoder) in GPT-3 175B and Bloom 176B and run our optimization algorithm to obtain the time–energy frontier. We perform strong scaling when varying the number of GPUs (Table 5) in order to keep the global batch size constant (largebatch1, ; largebatch2, ). We used A100 SXM GPUs for emulation, which we believe are more representative of large-scale training infrastructure.
@@
We compare the percentage of energy savings vs. running all maximum frequencies given by our emulator and real experiments for our A100 workloads and find that the emulator always underestimates energy savings. Specifically, savings on the leftmost and rightmost point of the frontier are underestimated by 18.6% and 21.7% on average, respectively. We believe this is due to our simplifying assumption that $P_{\mathrm{blocking}}$ is constant regardless of the GPU’s frequency. This means the savings given by the emulator can be considered a lower bound for actual savings.
@@
Result Summary.
Figure 7 breaks down the amount of energy bloat reduction for GPT-3 175B and Bloom 176B when slowdown degree is 1.2 on 1,024 GPUs. EnvPipe can only reduce intrinsic bloat as it does not provide a time–energy frontier; even for intrinsic bloat, it is suboptimal. In contrast, Perseus reduces energy consumption by up to 30% by reducing both intrinsic and extrinsic energy bloat.
@@
Table 6 shows Perseus’s intrinsic energy bloat reduction without stragglers for GPT-3 175B and Bloom 176B. The number of microbatches is varied based on Table 5. For all models, as more microbatches are added to the pipeline, the amount of intrinsic bloat decreases. This is fundamentally due to the ratio of microbatches in the 1F1B’s warm-up and flush phase (beginning and end) versus the steady state phase (middle). Most microbatches in the warm-up and flush phases can slow down until their minimum energy frequency, yielding large energy savings. However, microbatches in the pipeline’s steady state cannot slow down to their full potential when the amount of stage imbalance is not large, thereby yielding modest savings. When the number of microbatches in the pipeline increases, only the number of steady state microbatches increases, and energy reduction converges to the average energy savings of steady state microbatches.
@@
We introduce stragglers of varying slowdowns in large-scale emulation. Figure 8 reports the amount of intrinsic + extrinsic energy bloat reduction achieved by Perseus. The trend where energy saving increases until $T^{\prime}<T^{*}$ and wanes afterward is consistent with Section 6.2.2.
An interesting observation here is that there is a tradeoff between scale and energy savings: configurations with more pipelines have less percentage of energy savings or less amount of energy savings per pipeline. It may seem intuitive to assume that more pipelines bring more energy savings, as there is only one straggler pipeline and all the other pipelines can reduce their energy consumption. However, this holds only in weak scaling scenarios, where the per-pipeline batch size is held constant (i.e., increasing the global batch size proportionally with the number of pipelines). Instead, in the more realistic strong scaling configuration (i.e., the global batch size is constant and per-pipeline batch size is decreased as more pipelines deployed), each pipeline’s number of microbatches changes. With fewer microbatches, the ratio of pipeline bubble (time that GPUs are idle) at the beginning and end of each pipeline iteration increases (megatronlm-sc21, ). These bubbles cannot be perfectly eliminated by intrinsic energy bloat reduction, resulting in a smaller energy savings percentages. However, as the absolute number of GPUs increases, even a small savings percentage is expected to yield huge absolute energy savings.
@@ 6.4. Iteration Time–Energy Frontier Comparison
The energy bloat reductions in Sections 6.2 and 6.3 were made possible by the time–energy frontier obtained using Perseus’s optimization algorithm (§4). Here, we further examine the frontier with different parallelization configurations and models and compare against Zeus (zeus-nsdi23, ). Since Zeus only produces the training time–energy frontier for single-GPU training jobs, we implemented two Zeus-based baselines for large model training scenarios:
@@
Figure 9 shows the frontiers of all solutions for different sizes of GPT-3 under three parallelization configurations: (a) four stage pipeline parallelism on A100; (b) eight stage pipeline parallelism on A40; and (c) 3D parallelism (data parallelism 2, tensor parallelism 2, pipeline parallelism 4) on A40. All results were obtained from running on testbed GPUs. Appendix H has frontiers for other models.
Perseus Pareto-dominates both baselines derived from Zeus. ZeusGlobal is unaware of pipeline stage imbalances and slows down every stage, and therefore is unable to intrinsic energy bloat. ZeusPerStage can balance the forward computation time of each stage, but is unaware of the critical path of the DAG, slowing down critical computations. In contrast, Perseus can precisely slow down non-critical computations, effectively reducing energy bloat.
@@ 6.5. Overhead of Perseus
At the beginning of training, the client-side profiler (§5) profiles forward and backward computations in each stage. For our A100 workloads, the initial profiling phase increased end-to-end training time by 13 minutes on average, which is negligible overhead for large model training.
@@
The average runtime of the optimization algorithm (§4) across our A100 workloads was 6.5 minutes, with the longest being Bloom 3B (15.7 minutes). For our largest-scale emulation workload (GPT-3 175B on A100 with 8,192 GPUs), the algorithm took 87 seconds, which is short because it is sufficient to optimize just one data parallel pipeline (§4.4). While runtime will increase with larger DAGs for larger models, we believe the overhead is justified because training time is likely to also increase with the scale of the training job. Looking up the optimal energy schedule given the straggler’s iteration time $T^{\prime}$ is instantaneous.
7. Related Work
@@
Many recent works focus on enabling and accelerating large model training using 3D parallelism. GPipe (gpipe-neurips19, ) and PipeDream (pipedream-sosp19, ) first introduced pipeline parallelism. 3D parallelism, especially popularized for Transformer-based models by Megatron-LM (megatronlm-arxiv, ; megatronlm-sc21, ), is considered to be the go-to solution for modern large model training due to strong open-source projects (megatronlm-github, ; deepspeed-github, ; gpt-neox-github, ) and relatively low implementation complexity. Extending this, Alpa (alpa-osdi22, ), GSPMD (spmd_socc22, ), and nnScaler (nnscaler-osdi24, ) provide automatic parallelization for general DNNs. However, energy consumption is not an optimization metric for any of the major large model training frameworks.
Several works utilize computation idle times within large model training pipelines to insert additional useful work (bamboo-nsdi23, ; pipefisher-mlsys23, ; hydro-osdi23, ). In contrast, Perseus’s approach is to slow down preceding computations to fill the idle time and reduce power and energy consumption. The rationale is twofold. First, the amount of idle time–especially in the steady state of the pipeline–is typically short. Inserting extra work may slow down the entire pipeline or require pipeline re-partitioning to make it feasible. Second, GPUs, especially those in the earlier stage of the pipeline, are likely already constrained by memory capacity, leaving little headroom for extra computation. We note that even with extra computation inserted, some idle time is likely to remain, leaving room for Perseus to provide energy savings.
Deep Learning and Energy Consumption.
A line of work measures or estimates the large amount of energy consumption and carbon emission of deep learning workloads (patterson2021carbon, ; energy-nlp-policy, ; quantify-carbon, ; dodge-faact22, ; bloomcarbon-arxiv23, ). In terms of optimization, some works determine the GPU’s execution speed for a single fixed sequence of GPU computations (odpp-ccgrid20, ; datadriven-ccgrid20, ; gpoeo-tpds21, ; chase-ccai23, ; depo-future23, ; cofris-igsc23, ), falling short when the time–energy frontier of a complex large model computation DAG needs to be characterized. Zeus (zeus-nsdi23, ) is a recent work that observes the tradeoff between GPU computation time and energy consumption, but still focuses on single-GPU training. EnvPipe (envpipe-atc23, ), on the other hand, aims to find a point solution that reduces the energy consumption of large model training with minimum slowdown. However, EnvPipe’s heuristic assumes that the last pipeline stage is always the bottleneck, leading to suboptimal savings. In terms of optimizing a single training job, Perseus is a superset of both Zeus and EnvPipe, achieved by viewing large model training as a computation DAG and introducing a principled optimization algorithm.
Big Data and Energy Consumption.
There are existing works that change the frequency of CPUs to improve the energy efficiency of big data workloads, including those that only execute computation-intensive phases (e.g., map and reduce) with the maximum frequency (mapreduce-dvfs-igcc11, ; hadoop-dvfs-fgcs16, ), those that choose the lowest frequency that can meet a given deadline (hadoop-dvfs-js17, ; express-icac17, ), or those that choose frequencies that balance the completion time of parallel tasks (yet simply minimizing idle time for smaller tasks within one stage) (spark-dvfs-js21, ). Perseus provides a principled approach to setting the job’s end-to-end completion time (§3.1) and task execution speed (§4), and can be directly applied to big data workloads by generating a DAG of CPU computations and optimizing the CPU’s frequencies.
@@ 8. Conclusion
We presented Perseus, a software-based energy optimization system for large model training. Perseus builds on the observation that there are computation imbalances at different levels in large model training that cause intrinsic and extrinsic energy bloat, and simultaneously reduces both with a principled graph cut-based algorithm. As a result, Perseus advances the state-of-the-art of DNN training energy optimization by delivering energy savings with negligible slowdown, thereby also reducing average power draw and enhancing the sustainability of large model training.
Acknowledgements
We would like to thank the SOSP reviewers, our shepherd Eddie Kohler, and SymbioticLab members for their insightful feedback. This work is in part supported by NSF grants CNS-2104243, CNS-2106184, and CCF-2327011, NWO VICI grant 639.023.812, grants from VMware and the Mozilla Foundation, and a gift from Salesforce. We also thank Chameleon Cloud for providing A100 nodes as well as CloudLab. Jae-Won Chung is additionally supported by the Kwanjeong Educational Foundation.
@@ B.1. Minimum Imbalance Pipeline Partitioning
B.2. Does Imbalance Decrease with Larger Models?
Specifically for Transformer [77]-based models with homogeneous intermediate Transformer layers, the degree of imbalance will decrease if the number of pipeline stages is held constant. This is because the relative size of the embedding layer and language model head will decrease as intermediate Transformer layers increase in number and size. However, as the model grow larger, the number of pipeline stages will have to be increased simultaneously. This increases imbalance, because the relative amount of computation in one layer with respect to the amount in one pipeline stage becomes larger. As such, there is no simple relationship between model size and the amount of imbalance.
B.3. Alternative Planning Methods